OpenStudy (anonymous):

integral x^2 / (3x^2 +1)

myininaya (myininaya):

Have you tried trig sub?

OpenStudy (anonymous):

first you have to divide the numerator by denominator and then integrate.

OpenStudy (anonymous):

i tried but it didn't solve it

OpenStudy (loser66):

show your work, please, follow surjithayer's instruction.

hartnn (hartnn):

adjusting the numerator, x^2 = (1/3)(3x^2) = (1/3) [ (3x^2 +1) -1 ] then you can separate out the numerator and denominator...

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }\int\limits\limits_{}^{} \frac{ {[3x^{2}+1}]-1 }{ 3x ^{2}+1 }dx\] then \[\frac{ 1 }{ 3 } \int\limits_{}^{}1dx -\frac{ 1 }{ 3 }\int\limits_{}^{}\frac{ 1 }{ 3x ^{2}+1 }\] if it right then what i should do in the second part \[\frac{ 1 }{ 3 }\int\limits_{}^{}\frac{ 1 }{ 3x ^{2}+1 }dx\] because i stop here

hartnn (hartnn):

correct till there, now bring the denominator in the form x^2+a^2 like 3x^2+1 = (1/3) [x^2 + 1/3] = (1/3) [x^2 + (1/sqrt 3)^2] since we have the form, x^2+a^2, we can put x= a tan t here, x = tan t/ sqrt 3

OpenStudy (anonymous):

look what if we from the step let u=3x^2+1 then x^2=(u-1)/3 ok i think it easier

hartnn (hartnn):

letting u=3x^2+1 won't help. as du will be 6x dx and you don't have 'x' in numerator....

OpenStudy (anonymous):

when we let u the euqation will be \[\frac{ \frac{ u-1 }{ 3 } }{ u }\] then \[\frac{ { u-1 } }{ 3u }\] then integrate

hartnn (hartnn):

you still have dx part left to calculate...

hartnn (hartnn):

u-1/3u dx <<<<<this dx needs to be calculated in terms of du

hartnn (hartnn):

u=3x^2+1 du = 6x dx dx = du/6x = du/ (6\sqrt (u-1)/3) which makes it much more complicated...

hartnn (hartnn):

splitting and making the trigonometric substitution is the easiest way for this...

OpenStudy (anonymous):

thats right my mistake thank you

hartnn (hartnn):

no problem, ask me if you get stuck while doing trig. substitution.....or need any kind of help with this :)

hartnn (hartnn):

oh and since you're new here, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

thank you